Two inlet pipes can fill an empty swimming pool in 3 days and 6 days, respectively. A pump can empty a full pool in 4 days. How long will it take to fill the empty pool if both pipes are open and the pump is on?
Answer by richwmiller(17219)

1/x+1/y=1/p
1/3+1/6=1/p
2/6+1/6=1/p
3/6=1/p
1/2=1/p
p=2 days with just the inlets open
1/3+1/6-1/4=1/p
4/12+2/12-3/12=1/p
3/12=1/p
1/4=1/p
p=4 days with the inlets and the pump open